Announcements

The perfect balance of Reliability, Uptime & Brutal Performance!

[SOLVED] Orlando Outage

Sunday, July 2, 2017


[Sunday, 02 July 2017, 00:12]:  We are aware of the downage in Orlando. Our upstream are having connectivity issued. Expect further updates.




[solved] la3 emergency maintenance

Friday, February 10, 2017


10 February 2017, 11:38 - we are shutting down the LA3 server to replace a disk within. We do not expect this operation to take more than a few minutes. This announcement will be updated when the maintenance is complete.  10 February 2017, 13:33 - server is back online and containers are booted. If you're experiencing an issue with your VM, please open a ticket. Thank you. 

More »




[SOLVED] or5 node downtime

Monday, January 23, 2017


We are currently working on or5 to resolve the issue. Update 05:28 AM: Unfortunately, we do not have an update at this time. The datacenter staff is currently working on the issue. Update 05:36 AM: The datacenter staff has notified us that there is an issue with the memory modules of the server and that they are working on resolving it. We will update you once we have more information....

More »




[SOLVED] Orlando Outage

Thursday, December 1, 2016


 01 December 2016, 01:31 -  We are investigating an outage in our services in Orlando.  01 December 2016, 01:54 - There were some intermitent network issues. All services should be accessible now. 01 December 2016, 17:56 - this appears to have reoccurred. We have engaged our datacenter to investigate further.  01 December 2016, 18:14 - This issue is rectified and servers are back...

More »




[SOLVED] LA network outage

Friday, November 25, 2016


25 November 2016, 23:10 - We're investigating an outage of our LA services.  25 November 2016, 23:20 - Servers are back online.  25 November 2016, 23:34 - The datacenter has informed us they are investigating the reason for the outage, as it is not immediately apparent at this time.  




[SOLVED] Outage of uk2 and uk3

Thursday, September 22, 2016


 22 September 2016  |  15:01 PM - We're investigating a network connectivity issue of our uk2 and uk3 servers. We'll update this announcement as we have more information.   22 September 2016  |  15:11 PM - There is an issue with the BGP announcement of the IPs of these servers. We have escalated this with the datacenter at high priority.  22 September 2016  |  16:05 PM -...

More »




[SOLVED] Network instability in Orlando

Tuesday, September 13, 2016


09/13/2016 23:16 - We are acknowledging network instability in our Orlando location and have three network techs on the situation now. This may cause your VPS to appear "offline" in the client area. Please refrain from clicking start/stop/reset buttons randomly to "fix" the issue, as this will not fix the issue. We'll update this announcement as we have more information. 16 September 2016 ...

More »




[SOLVED] uk1 Emergency Maintenance

Monday, August 22, 2016


We are performing an emergency maintenance on uk1 node right now. Your VPS should be online in the next 15 minutes. Update 16:32: All VMs are back online again.




[SOLVED] Los Angeles Migration

Friday, April 8, 2016


Update 02:52: The migration is completed. We have just suspended your VPS's memory to disk and then restarted them. So, your VPS's didn't get rebooted and they're continuing working where they left off. We are aware that your VPS's in Los Angeles are down at the moment. As scheduled, Los Angeles physical data center migration is in progress now. We kindly ask you not to open a support...

More »




[SOLVED] UK Network Issues

Thursday, December 10, 2015


We are in contact with the DC about this for the last 3 days as it started happening. We are not getting much resolution. Here's best responses we got so far: "I apologize for the delay. We have been experiencing a large sophisticated Distributed Denial of Service attack against IPs from within the UK. We are monitoring this and working to mitigate it." "As far as we can tell it appears...

More »





« Prev Page
Next Page »

RSS View RSS Feed